# The Complexity of Sequential Prediction in Dynamical Systems

### v1

@@ G.2 Linear Boolean Networks
$\texttt{R}^{\texttt{MK}}_{\mathcal{A}}(T,\mathcal{F})\geq\frac{1}{2}\mathbb{E}\left[\left|\sum_{t=1}^{T}\sigma_{t}\mathbbm{1}\{x_{t-1}\notin S\}\right|\right]=\frac{1}{2}\mathbb{E}[|Y|]=\frac{\sqrt{T}}{16\sqrt{3}}.$
This completes our proof. ∎
Appendix I Tightness of Upperbound in Theorem 4.1
@@
Consider =$\mathcal{F}=\{f_{\sigma}\,:\,\sigma\in\{-1,1\}^{d}\}$. It is not too hard to see that $\texttt{L}(\mathcal{F})=d$. The fact that $\texttt{L}(\mathcal{F})\leq d$ is trivial because any Littlesone tree can only have $\{1,2,\ldots,d\}$ in the internal nodes. This is because all functions output the same state in the domain $\mathcal{X}\backslash[d]$. Since the internal nodes in any given path of the Littlestone tree need to be different, the depth of any shattered tree is $\leq d$. To see why $\texttt{L}(\mathcal{F})\geq d$, consider a complete binary tree $\mathcal{T}$ of depth $d$ with all the internal nodes in level $i\in[d]$ containing the element $i$. Let $-(d+i)$ and $(d+i)$ label the left and right outgoing edges respectively of every node in level $i$. Note that $\mathcal{T}$ is shattered by $\mathcal{F}$ as for any path $\varepsilon\in\{-1,1\}^{d}$ down the tree $\mathcal{T}$, there exists a $f_{\varepsilon}\in\mathcal{F}$ such that $f_{\varepsilon}(i)=\varepsilon_{i}(d+i)$ for all $i\in[d]$. Thus, we have shown that $\texttt{L}(\mathcal{F})=d$.
We now show that $\inf_{\mathcal{A}}\texttt{R}^{\texttt{MK}}_{\mathcal{A}}(T,\mathcal{F})=\Omega(\sqrt{T\texttt{L}(\mathcal{F})})$, proving that the upperbound in Theorem 4.1 is tight up to $\sqrt{\log{T}}$. For a $k\in\mathbb{N}$, pick $T=2kd-1$. Draw $\varepsilon\in\{-1,+1\}^{T}$ where $\varepsilon_{i}\sim\text{Uniform}(\{-1,1\})$ and consider a stream
@@
$\begin{split}\texttt{R}^{\texttt{MK}}_{\mathcal{A}}(T,\mathcal{F})&=\mathbb{E}\left[\sum_{t=1}^{T}\mathbbm{1}\{\mathcal{A}(x_{<t})\neq x_{t}\}\right]-\mathbb{E}\left[\inf_{f\in\mathcal{F}}\sum_{t=1}^{T}\mathbbm{1}\{f(x_{t-1})\neq x_{t}\}\right]\\ &\geq\frac{1}{2}\mathbb{E}\left[\sum_{i=1}^{d}\left|\sum_{t=2k(i-1)+1}^{2ki-1}\varepsilon_{t}\mathbbm{1}\{x_{t-1}=i\}\right|\right]-(d-1)\\ &\geq\frac{1}{2}\sum_{i=1}^{d}\sqrt{\frac{|\{2k(i-1)+1\leq t\leq 2ki-1\,:\,x_{t-1}=i\}|}{2}}-(d-1)\\ &=\frac{1}{2}\sum_{i=1}^{d}\sqrt{\frac{k}{2}}-(d-1).\end{split}$
where the second inequality follows due to Khinchine’s inequality [Cesa-Bianchi and Lugosi, 2006, Lemma 8.2]. The final equality holds because in each block of $2k$ rounds from $t=2k(i-1)+1$ to $t=2ki-1$, we have $x_{t-1}=i$ in exactly $k$ of those rounds. Using $T=2kd-1$, we further obtain $\texttt{R}^{\texttt{MK}}_{\mathcal{A}}(T,\mathcal{F})\geq\frac{1}{4}\sqrt{(T+1)d}-(d-1)=\Omega(\sqrt{Td})$ for $T\gg d$. This establishes our claim that the upperbound in Theorem 4.1 is tight up to $\sqrt{\log{T}}$.
Appendix J Proof of Theorem 5.1
@@
Proof.
(of the upperbound in Theorem 5.1) Let $\{x_{t}\}_{t=0}^{T}$ be the trajectory to be observed by the learner and $f^{\star}\in\operatorname*{arg\,min}_{f\in\mathcal{F}}\sum_{t=1}^{T}\mathbbm{1}\{x_{t}\neq f^{t}(x_{0})\}$ the optimal evolution function in hindsight. Given the time horizon $T$, let $L_{T}=\{L\subset[T]:|L|\leq\texttt{C}_{T}(\mathcal{F})\}$ denote the set of all possible subsets of $[T]$ with size at most $\texttt{C}_{T}(\mathcal{F})$. For every $L\in L_{T}$, let $\phi:L\rightarrow[K]$ denote a function mapping time points in $L$ to an integer in $[K]$. On time point $t\in L$, we should think of $\phi(t)$ as an index into the set $\mathcal{F}(x_{t-1})$. To that end, for an index $i\in[|\mathcal{F}(x)|]$, let $\mathcal{F}_{i}(x)$ denote the $i$-th element of the list obtained after sorting $\mathcal{F}(x)$ in its natural order. Let $\Phi_{L}=[K]^{L}$ denote all such functions $\phi$. For each $L\in L_{T}$ and $\phi\in\Phi_{L}$, define an expert $\mathcal{E}_{L,\phi}$. On time point $t\in[T]$, the prediction of expert $\mathcal{E}_{L,\phi}$ is defined recursively by
@@
We claim that there exists an expert $\mathcal{E}_{L^{\star},\phi^{\star}}$ such that $\mathcal{E}_{L^{\star},\phi^{\star}}(x_{0},t)=f^{\star,t}(x_{0})$ for all $t\in[T]$. To see this, consider the hypothetical trajectory $S^{\star}=\{f^{\star,t}(x_{0})\}_{t=1}^{T}$ generated by $f^{\star}$. Let $L^{\star}=\{t_{1},t_{2},...\}$ be the indices on which Algorithm 1 would have made a mistake had it run and updated on $S^{\star}$. By the guarantee of Algorithm 1, we know that $|L^{\star}|\leq\texttt{C}_{T}(\mathcal{F})$. Moreover, there exists a $\phi^{\star}\in\Phi_{L^{\star}}$ such that $\mathcal{F}_{\phi^{\star}(i)}(f^{\star,i-1}(x_{0}))=f^{\star,i}(x_{0})$ for all $i\in L^{\star}$. By construction of $\mathcal{E}$, there exists an expert $\mathcal{E}_{L^{\star},\phi^{\star}}$ parameterized by $L^{\star}$ and $\phi^{\star}$. We claim that $\mathcal{E}_{L^{\star},\phi^{\star}}(x_{0},t)=f^{\star,t}(x_{0})$ for all $t\in[T]$. This follows by strong induction on $t\in[T]$. For the base case $t=1$, there are two subcases to consider. If $1\in L^{\star}$, then we have that $\mathcal{E}_{L^{\star},\phi^{\star}}(x_{0},1)=\mathcal{F}_{\phi^{\star}(1)}(\mathcal{E}_{L^{\star},\phi^{\star}}(x_{0},0))=\mathcal{F}_{\phi^{\star}(1)}(x_{0})=f^{\star,1}(x_{0})$. If $1\notin L^{\star}$, then $\mathcal{E}_{L^{\star},\phi^{\star}}(x_{0},1)=\mathcal{A}\left(x_{0},1|\{\}\right)=f^{\star,1}(x_{0})$, where the last equality follows by definition of $L^{\star}$. Now for the induction step, suppose that $\mathcal{E}_{L^{\star},\phi^{\star}}(x_{0},i)=f^{\star,i}(x_{0})$ for all $i\leq t$. Then, if $t+1\in L^{\star}$, we have that $\mathcal{E}_{L^{\star},\phi^{\star}}(x_{0},t+1)=\mathcal{F}_{\phi^{\star}(t+1)}(\mathcal{E}_{L^{\star},\phi^{\star}}(x_{0},t))=\mathcal{F}_{\phi^{\star}(t+1)}(f^{\star,t}(x_{0}))=f^{\star,t+1}(x_{0}).$ If $t+1\notin L^{\star}$, then $\mathcal{E}_{L^{\star},\phi^{\star}}(x_{0},t+1)=\mathcal{A}\left(x_{0},t+1|\{\mathcal{E}_{L^{\star},\phi^{\star}}(x_{0},i)\}_{i=1}^{t}\right)=\mathcal{A}\left(x_{0},t+1|\{f^{\star,i}(x_{0})\}_{i=1}^{t}\right)=f^{\star,t+1}(x_{0})$, where the last equality again is due to the definition of $L^{\star}$.
Now, consider the learner that runs the celebrated Randomized Exponential Weights Algorithm, denoted hereinafter by $\mathcal{P}$, using the set of experts $E$ with learning rate $\eta=\sqrt{\frac{2\ln(|E|)}{T}}$. By Theorem 21.11 of Shalev-Shwartz and Ben-David [2014], we have that
@@
$\leq\inf_{\mathcal{E}\in E}\sum_{t=1}^{T}\mathbbm{1}\{\mathcal{E}(x_{0},t)\neq x_{t}\}+\sqrt{2\texttt{C}_{T}(\mathcal{F})\,T\ln\Bigl{(}\frac{KT}{\texttt{C}_{T}(\mathcal{F})}\Bigl{)}}$
$\leq\sum_{t=1}^{T}\mathbbm{1}\{\mathcal{E}_{L^{\star},\phi^{\star}}(x_{0},t)\neq x_{t}\}+\sqrt{2\texttt{C}_{T}(\mathcal{F})\,T\ln\Bigl{(}\frac{KT}{\texttt{C}_{T}(\mathcal{F})}\Bigl{)}}$
$=\sum_{t=1}^{T}\mathbbm{1}\{f^{\star,t}(x_{0})\neq x_{t}\}+\sqrt{2\texttt{C}_{T}(\mathcal{F})\,T\ln\Bigl{(}\frac{KT}{\texttt{C}_{T}(\mathcal{F})}\Bigl{)}}$
where the second inequality follows because $|E|=\sum_{i=0}^{\texttt{C}_{T}(\mathcal{F})}K^{i}\,\binom{T}{i}\leq\Bigl{(}\frac{KT}{\texttt{C}_{T}(\mathcal{F})}\Bigl{)}^{\texttt{C}_{T}(\mathcal{F})}$. This completes the proof as $\mathcal{P}$ achieves the stated upperbound on expected regret. ∎
@@
J.1 Tightness of Lowerbound in Theorem 5.1
@@
and consider the class =$\mathcal{F}=\Bigl{\{}f_{\sigma}:\sigma\in\{-1,1\}^{\mathbb{N}\cup\,\{0\}}\Bigl{\}}.$ By construction of $\mathcal{F}$, branching only occurs on states in $\{0,1,...,p-1\}$ and their negation. Moreover, given any initial state $x_{0}\in\mathcal{X}$ and a time horizon $T\in\mathbb{N}$, the trajectory of any evolution in $\mathcal{F}$ is some signed version of the sequence $|x_{0}|+1,...,|x_{0}|+T$. From Theorem 3.3, its not too hard to see that $\texttt{C}_{T}(\mathcal{F})=\min\{p,T\}$. We now show that $\inf_{\mathcal{A}}\texttt{R}^{\texttt{FL}}_{\mathcal{A}}(T,\mathcal{F})\leq\frac{\texttt{C}_{T}(\mathcal{F})}{2}$. Consider the learner $\mathcal{A}$ which, given initial state $x_{0}$, checks whether $|x_{0}|\leq p-1$. If $|x_{0}|\leq p-1$, the learner $\mathcal{A}$ samples a random sequence $\varepsilon\sim\{-1,1\}^{p-|x_{0}|}$ and plays $\varepsilon_{t}(|x_{0}|+t)$ for $t\leq p-|x_{0}|$ and $|x_{0}|+t$ in all future rounds $t>p-|x_{0}|$. If $|x_{0}|>p-1$, the learner $\mathcal{A}$ plays $|x_{0}|+t$ for all $t\geq 1$.
@@
$\texttt{R}^{\texttt{FL}}_{\mathcal{A}}(T,\mathcal{F})$ $=\mathbbm{E}\left[\sum_{t=1}^{T}\mathbbm{1}\{\mathcal{A}(x_{<t})\neq x_{t}\}-\inf_{f\in\mathcal{F}}\sum_{t=1}^{T}\mathbbm{1}\{f^{t}(x_{0})\neq x_{t}\}\right]$
@@
$\texttt{R}^{\texttt{FL}}_{\mathcal{A}}(T,\mathcal{F})$ $=\mathbbm{E}\left[\sum_{t=1}^{T}\mathbbm{1}\{\mathcal{A}(x_{<t})\neq x_{t}\}-\inf_{f\in\mathcal{F}}\sum_{t=1}^{T}\mathbbm{1}\{f^{t}(x_{0})\neq x_{t}\}\right]$
@@
where in the last equality we used the fact that $\varepsilon_{t}\sim\text{Unif}(-1,1)$. Thus, the expected regret of such an learner $\mathcal{A}$ is at most $\frac{\min\{\max\{p-|x_{0}|,0\},T\}}{2}$. Taking $x_{0}=0$, gives that $\inf_{\mathcal{A}}\texttt{R}^{\texttt{FL}}_{\mathcal{A}}(T,\mathcal{F})\leq\frac{\texttt{C}_{T}(\mathcal{F})}{2}.$
J.2 Tightness of Upperbound in Theorem 5.1
@@
Given any initial state $x_{0}\in\mathcal{X}$ and time horizon $T\in\mathbb{N}$, the trajectory $\{f^{t}_{\sigma,S}(x_{0})\}_{t=1}^{T}$ is some signed version of the sequence $(|x_{0}|+1,...,|x_{0}|+T)$, where the signs depend on $S$ and $\sigma$. More importantly, the trajectory $\{f^{t}_{\sigma,S}(x_{0})\}_{t=1}^{T}$ switches signs at most $p+1$ times. Finally, consider the evolution class =$\mathcal{F}=\{f_{\sigma,S}:\sigma\in\{-1,1\}^{\mathbb{N}\cup\{0\}},S\in\mathcal{G}\}.$ Since the trajectory of any evolution $f\in\mathcal{F}$ can switch signs at most $p+1$ times, it follows that $\texttt{C}_{T}(\mathcal{F})\leq p+1$. Moreover, by considering a trajectory tree with the root node labeled by 00 and the set of evolutions parameterized by the tuple $(0,1,...,p)\in\mathcal{G}$, its not hard to see that $\texttt{C}_{T}(\mathcal{F})\geq p+1.$ Thus, $\texttt{C}_{T}(\mathcal{F})=p+1$.
We now claim that $\inf_{\mathcal{A}}\texttt{R}^{\emph{{FL}}}_{\mathcal{A}}(T,\mathcal{F})\geq\sqrt{\frac{\texttt{C}_{T}(\mathcal{F})\,T}{8}}$, which shows that the upperbound in Theorem 5.1 is tight up to a logarithmic factor in $T$ since $\sup_{x\in\mathcal{X}}|\mathcal{F}(x)|=2$. Consider $T=k(p+1)$ for some odd $k\in\mathbb{N}$. For $\varepsilon\in\{-1,1\}^{T}$, define $\tilde{\varepsilon}_{i}=\text{sign}\Bigl{(}\sum_{t=(i-1)k+1}^{ik}\varepsilon_{t}\Bigl{)}$ for all $i\in\{1,2,...,p+1\}$. The game proceeds as follows. The adversary samples a string $\varepsilon\in\{-1,1\}^{T}$ uniformly at random and constructs the random trajectory $\{x_{t}\}_{t=0}^{T}$ where $x_{0}=0$ is the initial state and $x_{t}=\varepsilon_{t}t$ for all $t\geq 1$. The adversary then passes $\{x_{t}\}_{t=0}^{T}$ to the learner.
@@
where the final step follows upon using Khinchine’s inequality [Cesa-Bianchi and Lugosi, 2006]. Combining these two bounds above, we obtain,
@@
$\geq(p+1)\sqrt{\frac{k}{8}}=\sqrt{\frac{\texttt{C}_{T}(\mathcal{F})T}{8}},$
@@
Appendix K Proof of Lemma 5.2
@@
Therefore, we have $\texttt{R}^{\texttt{FL}}_{\mathcal{A}}(T,\mathcal{F})\geq\frac{T}{2}-\frac{T}{3}\geq\frac{T}{6}.$

### v2

@@ Appendix E Proofs for Linear Systems
$\operatorname{MR}_{\mathcal{A}}(T,\mathcal{F})\geq\frac{1}{2}\mathbb{E}\left[\left|\sum_{t=1}^{T}\sigma_{t}\mathbbm{1}\{x_{t-1}\notin S\}\right|\right]=\frac{1}{2}\mathbb{E}[|Y|]=\frac{\sqrt{T}}{16\sqrt{3}}.$
This completes our proof.
Appendix G Tightness of upper bound in Theorem 4.1
@@
Consider =$\mathcal{F}=\{f_{\sigma}\,:\,\sigma\in\{-1,1\}^{d}\}$. It is not too hard to see that $\operatorname{L}(\mathcal{F})=d$. The fact that $\operatorname{L}(\mathcal{F})\leq d$ is trivial because any Littlesone tree can only have $\{1,2,\ldots,d\}$ in the internal nodes. This is because all functions output the same state in the domain $\mathcal{X}\backslash[d]$. Since the internal nodes in any given path of the Littlestone tree need to be different, the depth of any shattered tree is $\leq d$. To see why $\operatorname{L}(\mathcal{F})\geq d$, consider a complete binary tree $\mathcal{T}$ of depth $d$ with all the internal nodes in level $i\in[d]$ containing the element $i$. Let $-(d+i)$ and $(d+i)$ label the left and right outgoing edges respectively of every node in level $i$. Note that $\mathcal{T}$ is shattered by $\mathcal{F}$ as for any path $\varepsilon\in\{-1,1\}^{d}$ down the tree $\mathcal{T}$, there exists a $f_{\varepsilon}\in\mathcal{F}$ such that $f_{\varepsilon}(i)=\varepsilon_{i}(d+i)$ for all $i\in[d]$. Thus, we have shown that $\operatorname{L}(\mathcal{F})=d$.
We now show that $\inf_{\mathcal{A}}\operatorname{MR}_{\mathcal{A}}(T,\mathcal{F})=\Omega(\sqrt{T\operatorname{L}(\mathcal{F})})$, proving that the upper bound in Theorem 4.1 is tight up to $\sqrt{\log{T}}$. For a $k\in\mathbb{N}$, pick $T=2kd-1$. Draw $\varepsilon\in\{-1,+1\}^{T}$ where $\varepsilon_{i}\sim\text{Uniform}(\{-1,1\})$ and consider a stream
@@
$\begin{split}\operatorname{MR}_{\mathcal{A}}(T,\mathcal{F})&=\mathbb{E}\left[\sum_{t=1}^{T}\mathbbm{1}\{\mathcal{A}(x_{<t})\neq x_{t}\}\right]-\mathbb{E}\left[\inf_{f\in\mathcal{F}}\sum_{t=1}^{T}\mathbbm{1}\{f(x_{t-1})\neq x_{t}\}\right]\\ &\geq\frac{1}{2}\mathbb{E}\left[\sum_{i=1}^{d}\left|\sum_{t=2k(i-1)+1}^{2ki-1}\varepsilon_{t}\mathbbm{1}\{x_{t-1}=i\}\right|\right]-(d-1)\\ &\geq\frac{1}{2}\sum_{i=1}^{d}\sqrt{\frac{|\{2k(i-1)+1\leq t\leq 2ki-1\,:\,x_{t-1}=i\}|}{2}}-(d-1)\\ &=\frac{1}{2}\sum_{i=1}^{d}\sqrt{\frac{k}{2}}-(d-1).\end{split}$
where the second inequality follows due to Khinchine’s inequality (Cesa-Bianchi and Lugosi, 2006, Lemma 8.2). The final equality holds because in each block of $2k$ rounds from $t=2k(i-1)+1$ to $t=2ki-1$, we have $x_{t-1}=i$ in exactly $k$ of those rounds. Using $T=2kd-1$, we further obtain $\operatorname{MR}_{\mathcal{A}}(T,\mathcal{F})\geq\frac{1}{4}\sqrt{(T+1)d}-(d-1)=\Omega(\sqrt{Td})$ for $T\gg d$. This establishes our claim that the upper bound in Theorem 4.1 is tight up to $\sqrt{\log{T}}$.
Appendix H Proof of Theorem 4.3
@@
Proof H.1.
(of the upper bound in Theorem 4.3) Let $\{x_{t}\}_{t=0}^{T}$ be the trajectory to be observed by the learner and $f^{\star}\in\operatorname*{arg\,min}_{f\in\mathcal{F}}\sum_{t=1}^{T}\mathbbm{1}\{x_{t}\neq f^{t}(x_{0})\}$ the optimal evolution function in hindsight. Given the time horizon $T$, let $L_{T}=\{L\subset[T]:|L|\leq\operatorname{C}_{T}(\mathcal{F})\}$ denote the set of all possible subsets of $[T]$ with size at most $\operatorname{C}_{T}(\mathcal{F})$. For every $L\in L_{T}$, let $\phi:L\rightarrow[K]$ denote a function mapping time points in $L$ to an integer in $[K]$. On time point $t\in L$, we should think of $\phi(t)$ as an index into the set $\mathcal{F}(x_{t-1})$. To that end, for an index $i\in[|\mathcal{F}(x)|]$, let $\mathcal{F}_{i}(x)$ denote the $i$-th element of the list obtained after sorting $\mathcal{F}(x)$ in its natural order. Let $\Phi_{L}=[K]^{L}$ denote all such functions $\phi$. For each $L\in L_{T}$ and $\phi\in\Phi_{L}$, define an expert $\mathcal{E}_{L,\phi}$. On time point $t\in[T]$, the prediction of expert $\mathcal{E}_{L,\phi}$ is defined recursively by
@@
We claim that there exists an expert $\mathcal{E}_{L^{\star},\phi^{\star}}$ such that $\mathcal{E}_{L^{\star},\phi^{\star}}(x_{0},t)=f^{\star,t}(x_{0})$ for all $t\in[T]$. To see this, consider the hypothetical trajectory $S^{\star}=\{f^{\star,t}(x_{0})\}_{t=1}^{T}$ generated by $f^{\star}$. Let $L^{\star}=\{t_{1},t_{2},...\}$ be the indices on which Algorithm 1 would have made a mistake had it run and updated on $S^{\star}$. By the guarantee of Algorithm 1, we know that $|L^{\star}|\leq\operatorname{C}_{T}(\mathcal{F})$. Moreover, there exists a $\phi^{\star}\in\Phi_{L^{\star}}$ such that $\mathcal{F}_{\phi^{\star}(i)}(f^{\star,i-1}(x_{0}))=f^{\star,i}(x_{0})$ for all $i\in L^{\star}$. By construction of $\mathcal{E}$, there exists an expert $\mathcal{E}_{L^{\star},\phi^{\star}}$ parameterized by $L^{\star}$ and $\phi^{\star}$. We claim that $\mathcal{E}_{L^{\star},\phi^{\star}}(x_{0},t)=f^{\star,t}(x_{0})$ for all $t\in[T]$. This follows by strong induction on $t\in[T]$. For the base case $t=1$, there are two subcases to consider. If $1\in L^{\star}$, then we have that $\mathcal{E}_{L^{\star},\phi^{\star}}(x_{0},1)=\mathcal{F}_{\phi^{\star}(1)}(\mathcal{E}_{L^{\star},\phi^{\star}}(x_{0},0))=\mathcal{F}_{\phi^{\star}(1)}(x_{0})=f^{\star,1}(x_{0})$. If $1\notin L^{\star}$, then $\mathcal{E}_{L^{\star},\phi^{\star}}(x_{0},1)=\mathcal{A}\left(x_{0},1|\{\}\right)=f^{\star,1}(x_{0})$, where the last equality follows by definition of $L^{\star}$. Now for the induction step, suppose that $\mathcal{E}_{L^{\star},\phi^{\star}}(x_{0},i)=f^{\star,i}(x_{0})$ for all $i\leq t$. Then, if $t+1\in L^{\star}$, we have that $\mathcal{E}_{L^{\star},\phi^{\star}}(x_{0},t+1)=\mathcal{F}_{\phi^{\star}(t+1)}(\mathcal{E}_{L^{\star},\phi^{\star}}(x_{0},t))=\mathcal{F}_{\phi^{\star}(t+1)}(f^{\star,t}(x_{0}))=f^{\star,t+1}(x_{0}).$ If $t+1\notin L^{\star}$, then $\mathcal{E}_{L^{\star},\phi^{\star}}(x_{0},t+1)=\mathcal{A}\left(x_{0},t+1|\{\mathcal{E}_{L^{\star},\phi^{\star}}(x_{0},i)\}_{i=1}^{t}\right)=\mathcal{A}\left(x_{0},t+1|\{f^{\star,i}(x_{0})\}_{i=1}^{t}\right)=f^{\star,t+1}(x_{0})$, where the last equality again is due to the definition of $L^{\star}$.
Now, consider the learner that runs the celebrated Randomized Exponential Weights Algorithm, denoted hereinafter by $\mathcal{P}$, using the set of experts $E$ with learning rate $\eta=\sqrt{\frac{2\ln(|E|)}{T}}$. By Theorem 21.11 of Shalev-Shwartz and Ben-David (2014), we have that
@@
$\leq\inf_{\mathcal{E}\in E}\sum_{t=1}^{T}\mathbbm{1}\{\mathcal{E}(x_{0},t)\neq x_{t}\}+\sqrt{2\operatorname{C}_{T}(\mathcal{F})\,T\ln\Bigl{(}\frac{KT}{\operatorname{C}_{T}(\mathcal{F})}\Bigl{)}}$
$\leq\sum_{t=1}^{T}\mathbbm{1}\{\mathcal{E}_{L^{\star},\phi^{\star}}(x_{0},t)\neq x_{t}\}+\sqrt{2\operatorname{C}_{T}(\mathcal{F})\,T\ln\Bigl{(}\frac{KT}{\operatorname{C}_{T}(\mathcal{F})}\Bigl{)}}$
$=\sum_{t=1}^{T}\mathbbm{1}\{f^{\star,t}(x_{0})\neq x_{t}\}+\sqrt{2\operatorname{C}_{T}(\mathcal{F})\,T\ln\Bigl{(}\frac{KT}{\operatorname{C}_{T}(\mathcal{F})}\Bigl{)}}$
where the second inequality follows because $|E|=\sum_{i=0}^{\operatorname{C}_{T}(\mathcal{F})}K^{i}\,\binom{T}{i}\leq\Bigl{(}\frac{KT}{\operatorname{C}_{T}(\mathcal{F})}\Bigl{)}^{\operatorname{C}_{T}(\mathcal{F})}$. This completes the proof as $\mathcal{P}$ achieves the stated upper bound on expected regret.
@@
H.1 Tightness of lower bound in Theorem 4.3
@@
and consider the class =$\mathcal{F}=\Bigl{\{}f_{\sigma}:\sigma\in\{-1,1\}^{\mathbb{N}\cup\,\{0\}}\Bigl{\}}.$ By construction of $\mathcal{F}$, branching only occurs on states in $\{0,1,...,p-1\}$ and their negation. Moreover, given any initial state $x_{0}\in\mathcal{X}$ and a time horizon $T\in\mathbb{N}$, the trajectory of any evolution in $\mathcal{F}$ is some signed version of the sequence $|x_{0}|+1,...,|x_{0}|+T$. From Theorem 3.2, its not too hard to see that $\operatorname{C}_{T}(\mathcal{F})=\min\{p,T\}$. We now show that $\inf_{\mathcal{A}}\operatorname{FR}_{\mathcal{A}}(T,\mathcal{F})\leq\frac{\operatorname{C}_{T}(\mathcal{F})}{2}$. Consider the learner $\mathcal{A}$ which, given initial state $x_{0}$, checks whether $|x_{0}|\leq p-1$. If $|x_{0}|\leq p-1$, the learner $\mathcal{A}$ samples a random sequence $\varepsilon\sim\{-1,1\}^{p-|x_{0}|}$ and plays $\varepsilon_{t}(|x_{0}|+t)$ for $t\leq p-|x_{0}|$ and $|x_{0}|+t$ in all future rounds $t>p-|x_{0}|$. If $|x_{0}|>p-1$, the learner $\mathcal{A}$ plays $|x_{0}|+t$ for all $t\geq 1$.
@@
$\operatorname{FR}_{\mathcal{A}}(T,\mathcal{F})$ $=\mathbbm{E}\left[\sum_{t=1}^{T}\mathbbm{1}\{\mathcal{A}(x_{<t})\neq x_{t}\}-\inf_{f\in\mathcal{F}}\sum_{t=1}^{T}\mathbbm{1}\{f^{t}(x_{0})\neq x_{t}\}\right]$
@@
$\operatorname{FR}_{\mathcal{A}}(T,\mathcal{F})$ $=\mathbbm{E}\left[\sum_{t=1}^{T}\mathbbm{1}\{\mathcal{A}(x_{<t})\neq x_{t}\}-\inf_{f\in\mathcal{F}}\sum_{t=1}^{T}\mathbbm{1}\{f^{t}(x_{0})\neq x_{t}\}\right]$
@@
where in the last equality we used the fact that $\varepsilon_{t}\sim\text{Unif}(-1,1)$. Thus, the expected regret of such an learner $\mathcal{A}$ is at most $\frac{\min\{\max\{p-|x_{0}|,0\},T\}}{2}$. Taking $x_{0}=0$, gives that $\inf_{\mathcal{A}}\operatorname{FR}_{\mathcal{A}}(T,\mathcal{F})\leq\frac{\operatorname{C}_{T}(\mathcal{F})}{2}.$
H.2 Tightness of upper bound in Theorem 4.3
@@
Given any initial state $x_{0}\in\mathcal{X}$ and time horizon $T\in\mathbb{N}$, the trajectory $\{f^{t}_{\sigma,S}(x_{0})\}_{t=1}^{T}$ is some signed version of the sequence $(|x_{0}|+1,...,|x_{0}|+T)$, where the signs depend on $S$ and $\sigma$. More importantly, the trajectory $\{f^{t}_{\sigma,S}(x_{0})\}_{t=1}^{T}$ switches signs at most $p+1$ times. Finally, consider the evolution class =$\mathcal{F}=\{f_{\sigma,S}:\sigma\in\{-1,1\}^{\mathbb{N}\cup\{0\}},S\in\mathcal{G}\}.$ Since the trajectory of any evolution $f\in\mathcal{F}$ can switch signs at most $p+1$ times, it follows that $\operatorname{C}_{T}(\mathcal{F})\leq p+1$. Moreover, by considering a trajectory tree with the root node labeled by 00 and the set of evolutions parameterized by the tuple $(0,1,...,p)\in\mathcal{G}$, its not hard to see that $\operatorname{C}_{T}(\mathcal{F})\geq p+1.$ Thus, $\operatorname{C}_{T}(\mathcal{F})=p+1$.
We now claim that $\inf_{\mathcal{A}}\operatorname{FR}_{\mathcal{A}}(T,\mathcal{F})\geq\sqrt{\frac{\operatorname{C}_{T}(\mathcal{F})\,T}{8}}$, which shows that the upper bound in Theorem 4.3 is tight up to a logarithmic factor in $T$ since $\sup_{x\in\mathcal{X}}|\mathcal{F}(x)|=2$. Consider $T=k(p+1)$ for some odd $k\in\mathbb{N}$. For $\varepsilon\in\{-1,1\}^{T}$, define $\tilde{\varepsilon}_{i}=\text{sign}\Bigl{(}\sum_{t=(i-1)k+1}^{ik}\varepsilon_{t}\Bigl{)}$ for all $i\in\{1,2,...,p+1\}$. The game proceeds as follows. The adversary samples a string $\varepsilon\in\{-1,1\}^{T}$ uniformly at random and constructs the random trajectory $\{x_{t}\}_{t=0}^{T}$ where $x_{0}=0$ is the initial state and $x_{t}=\varepsilon_{t}t$ for all $t\geq 1$. The adversary then passes $\{x_{t}\}_{t=0}^{T}$ to the learner.
@@
where the final step follows upon using Khinchine’s inequality (Cesa-Bianchi and Lugosi, 2006). Combining these two bounds above, we obtain,
@@
$\geq(p+1)\sqrt{\frac{k}{8}}=\sqrt{\frac{\operatorname{C}_{T}(\mathcal{F})T}{8}},$
@@
Appendix I Proof of Theorem 4.4
@@
Therefore, we have $\operatorname{FR}_{\mathcal{A}}(T,\mathcal{F})\geq\frac{T}{2}-\frac{T}{3}\geq\frac{T}{6}.$